# FAST-GSC: Fast and Adaptive Semantic Transmission for Generative Semantic Communication

### v1

@@ VI Conclusion
In this paper, we have proposed a fast and adaptive semantic transmission design to reduce task latency within the conventional GSC framework while maintaining high task performance. Specifically, to decrease task latency, we have designed a communication mechanism that incorporates parallel semantic extraction and inference. Based on the direct implementations of the proposed mechanism, we have observed that the sequential dependencies among various semantic units and the reception immediacy of individual units significantly influence its task performance. Given that, we have implemented a RL-guided temporal prompt engineering method at the transmitter to sequence the extraction models according to their semantic sequential dependencies. Furthermore, we have mitigated the stringent reception immediacy by proposing a sequential conditional denoising approach based on a semantic difference calculation module at the receiver. Simulation results have validated the effectiveness of our proposed framework and demonstrated its significant potential for maintaining high task performance while reducing task latency.
Looking to the future, our devised framework can be extended to multi-user communication scenarios. By segmenting large semantic information blocks into smaller units, we enable more fine-grained semantic importance analysis and adaptive transmission design according to the users’ diverse task requirements. This approach also reduces the data delivered to a user at any given time, thereby mitigating network congestion in multi-user communication networks. Besides, our framework can also be utilized to enhance the communication efficiency for multi-modal data transmission. Specifically, it can be adapted to dynamically determine the transmission priority and resource allocation for different types of data, such as text, images, and video, based on their semantic significance and the current network conditions. Such directions are critical for the development of GSC and are worth further research.

### v2

@@ VI Conclusion
This paper has introduced FAST, a flexible and adaptive semantic transmission framework designed to address the challenges of complex applications such as digital twin systems and virtual reality environments. To overcome the computational constraints in such environments, we have developed a novel sequential semantic extraction approach guided by temporal prompt engineering, enabling efficient distillation and transmission of semantic units. At the receiver side, a sequential conditional denoising module has been implemented to progressively reconstruct content using the streamed semantic information. For multi-user scenarios, we have created a semantic-aware resource allocation scheme that dynamically distributes bandwidth through joint optimization of semantic dependencies, task priorities, and channel states. Experimental results have confirmed that the proposed framework has maintained reconstruction precision comparable to conventional GSC systems while achieving substantial improvements in both latency reduction and overall system efficiency. These findings have demonstrated its enhanced potential for deployment in multi-user GSC scenarios characterized by stringent communication and computational constraints.
Looking forward, several research directions merit further investigation. First, extending the FAST framework to support more diverse generative models beyond diffusion models represents a promising avenue. Second, exploring cross-modal semantic alignment techniques could enhance the framework’s capability in handling increasingly heterogeneous data sources. Finally, investigating security and privacy preservation mechanisms for semantic transmission in sensitive applications remains an important challenge for future work.